A car with a mass 700 kg accelerates uniformly from rest to a velocity of m s-1 in a time of 10 s. the frictional resistance to the motion of the vehicle is assumed to be constant throughout at 0.5 kN. calculate the force accelerating the car.
1
Expert's answer
2020-04-10T13:40:16-0400

Assume that "v=30m\/s".


"ma=F-F_f" "\to" "F=ma+F_f"


"v=at"


"F=ma+F_f=m\frac{v}{t}+F_f=700\frac{30}{10}+500=2600N=2.6kN"